OpenStudy (anonymous):

A 15-foot ladder leans against a wall at a 60 degree angle of elevation. What is the distance from the ladder's base to the wall?

OpenStudy (anonymous):

Cos60= x/15 x=15cos60 x=15/2 or 7.5

OpenStudy (whpalmer4):

A picture is usually a good place to start: 

OpenStudy (whpalmer4):

Naturally, we have to assume that the angle between ground and wall is a right angle. After that, remembering that the \(\cos\) function is the fraction \(\large\frac{\text{adjacent}}{\text{hypotenuse}}\) leads to the result @MelissaVictoria4002 gave you.